# Adversarial Robustness for Deep Learning-based Wildfire Prediction Models

### v2

@@ 6 Conclusions
WARP supports both global and local adversarial attack methods. While the global attack method employs image-contextualized random noise overlays, the local attack method is tailored to address two key aspects of smoke detection: 1) classification between smoke-positive and smoke-negative instances and 2) smoke object bounding box localization.
In a comprehensive comparative study, we assessed the adversarial robustness of CNN- and Transformer-based real-time smoke detection models, leveraging WARP’s model-agnostic capabilities. The main findings are summarized below:
  • The Global Sanity Check revealed that the transformer-based RT-DETR-l model was substantially more vulnerable to global noise injection compared to the CNN-based YOLOv8, even at noise levels barely visible to the human eye.
  • Analysis of the Classification Flip Probabilities revealed that both CNN- and Transformer-based models are extremely sensitive to local noise injection. A single injected noise grid (out of 625 total) resulted in flip probabilities of $\approx 50\%$ and $\approx 34\%$ in the smoke-positive prediction for the CNN and Transformer models, respectively. These results underscore the need for further model training using data augmentation techniques.
  • Analysis of the Localization Deception Rate revealed areas within the images that were particularly vulnerable to local noise injection. Detailed analysis suggests that human annotation bias may contribute to this vulnerability in the middle-horizontal region, offering insights for future data augmentation strategies to enhance model robustness.
Based on these findings, we proposed wildfire-specific data augmentation approaches. We leave the detailed analysis of the proposed data augmentation approaches in future studies.
@@ A.1 Definition of mAP
Precision and Recall are well established performance metrics in machine learning. Average Precision can be calculated by taking the area beneath the Precision-Recall curve (i.e., $p(r)$), typically using the Trapezoid Rule.

### v3

@@ 6 Conclusions
WARP supports both global and local adversarial attack methods. While the global attack method employs image-contextualized random noise overlays, the local attack method is tailored to address two key aspects of smoke detection: (1) classification between smoke-positive and smoke-negative instances and (2) smoke object bounding box localization.
Leveraging WARP’s model agnostic capabilities, we trained and compared CNN- and transformer-based wildfire detection models. We found in the global attack method that transformers suffer from higher precision degradation under global noise. In the local attack method, we found that while CNNs’ classification results are less robust than those of transformers, CNNs performed better in terms of localization ability. Future studies regarding this perturbation method are recommended to obtain a conclusive result. Finally, an auxiliary test suggests that annotation bias may be partially responsible for deceiving both models with local noise patches. Based on these findings, we proposed wildfire-specific data augmentation approaches. We leave a detailed analysis of the proposed data augmentation approaches for future studies.
WARP is a large step towards making DNN-based wildfire detection models more practical. Full-scale implementation of these systems will inevitably be costly. With the already high firefighting costs in the US, vulnerabilities to even simple adversarial attacks add great uncertainty in the reliability of these models. By identifying these biases and creating simple countermeasures in the form of data augmentation, WARP ensures both model accountability and dependability. We hope that this work will advance wildfire detection models enough to fully implement a completely automatic prediction system against wildfires, which can save the valuable lives and infrastructure that are lost to it.
@@ A.1 Definition of mAP
Precision and Recall are well established performance metrics in machine learning. Average Precision can be calculated by taking the area beneath the Precision–Recall curve (i.e., $p(r)$), typically using the Trapezoid Rule.